OpenStudy (anonymous):

Solve the following system of equations: 1/2 x + 1/3y = 7 1/4x + 2/3y = 6 how do I solve it?

OpenStudy (anonymous):

Multiple choices: 1. (1/6,14) 2. (6,12) 3. (10 2/3,5) 4. (2 4/9, -1/6)

OpenStudy (anonymous):

Are the fractions confusing you? If so, just multiply the first equation by 6 and the second by 12. 3x + 2y = 42 3x + 8y = 72 Can you solve now?

OpenStudy (anonymous):

not really, what exactly am i suppose to multiply to get the answer?

OpenStudy (anonymous):

It's rather a process that just one step. Multiply the first equation by -1. Then add vertically combining like terms. -3x - 2y = -42 3x + 8y = 72 The x values cancel. 6y = 30 Solve for x here and plug hat in for y here and solve for x: 3x + 2y = 42

OpenStudy (anonymous):

I am sorry but i am so bad at math, I still have no idea what you're talking about...

OpenStudy (anonymous):

and how to do it

OpenStudy (anonymous):

Alright. THere are several methods for this, but the most basic which you probably should eb using are elimination, graphing and substitution. Here are a few links. The first is elimination and the second is substitution. The graphing part can be done by graphing the lines and seeing where they cross. http://www.purplemath.com/modules/systlin5.htm http://www.purplemath.com/modules/systlin4.htm

OpenStudy (anonymous):

If you still have questions, you can ask me here.

OpenStudy (anonymous):

3/4 + 0.9 =13?

OpenStudy (anonymous):

First of all, when you add them, it should be 3/4x + y = 13 because 1/3 + 2/3 = 1. Second, the purpose of elimination is to get one of the terms to cancel so that you can solve for a variable. If you have 3/4x + y = 13, then you can't solve for x or y can you since you don't know either value.

OpenStudy (anonymous):

If you go farther down the link, it'll show you what to do for that kind of thing.

OpenStudy (anonymous):

For the frist part in the first sentence, it should be 1/3x + 2/3y = 13

OpenStudy (anonymous):

wow, so confusing but okay so I got that the first equation is changed to 1/3x + 2/3y = 13

OpenStudy (anonymous):

Double mistake. When I put: First of all, when you add them, it should be 3/4x + y = 13 because 1/3 + 2/3 = 1. The "1/3 + 2/3 = 1" should be 3/4x + y = 13

OpenStudy (anonymous):

omg... okay, i am dying here ... 3/4x + y = 13 got it

OpenStudy (anonymous):

then what do i do next?

OpenStudy (anonymous):

I'll show you what I did to solve this using elimination and you can browse what they did for the other examples later. 1/2 x + 1/3y = 7 1/4x + 2/3y = 6 1/2 x + 1/3y = 7 -2(1/4x + 2/3y = 6) 1/2 x + 1/3y = 7 -1/2x - 4/3y = -12 ADD VERTICALLY! 1/2 x + 1/3y = 7 -1/2x - 4/3y = -12 ------------------ 0 - y = -5 -y = -5 y = 5 Then plug this in for y and solve for x. 1/2x + 1/3y = 7 1/2x + 1/3(5) = 7 1/2x + 5/3 = 7 1/2x = 7 - 5/3 1/2x = 16/3 x = 32/3 which when you put into mixed number form is x = 10 2/3 Then, (10 2/3, 5) is the answer. Do you get it?

OpenStudy (anonymous):

why did you use (-2)? 1/2 x + 1/3y = 7 -2(1/4x + 2/3y = 6)

OpenStudy (anonymous):

I used -2 because -2(1/4) = -1/2 right? See how if you add 1/2 and -1/2, they turn to 0?

OpenStudy (anonymous):

okay, thank you so much for your help and time.. :)

OpenStudy (anonymous):

Anytime :)